# Physics-informed DeepCT: Sinogram Wavelet Decomposition Meets Masked Diffusion

### v1

@@ Abstract
Diffusion model shows remarkable potential on sparse-view computed tomography (SVCT) reconstruction. However, when a network is trained on a limited sample space, its generalization capability may be constrained, which degrades performance on unfamiliar data. For image generation tasks, this can lead to issues such as blurry details and inconsistencies between regions. To alleviate this problem, we propose a Sinogram-based Wavelet random decomposition And Random mask diffusion Model (SWARM) for SVCT reconstruction. Specifically, introducing a random mask strategy in the sinogram effectively expands the limited training sample space. This enables the model to learn a broader range of data distributions, enhancing its understanding and generalization of data uncertainty. In addition, applying a random training strategy to the high-frequency components of the sinogram wavelet enhances feature representation and improves the ability to capture details in different frequency bands, thereby improving performance and robustness. Two-stage iterative reconstruction method is adopted to ensure the global consistency of the reconstructed image while refining its details. Experimental results demonstrate that SWARM outperforms competing approaches in both quantitative and qualitative performance across various datasets.
{IEEEkeywords}
@@ 1 Introduction
In recent years, diffusion models have been increasingly employed in SVCT reconstruction owing to their superior capabilities. Xia et al.[19] introduced a patch-based denoising probabilistic diffusion model to improve SVCT reconstruction performance and address large memory requirements. Additionally, Wu et al.[20] proposed an iteratively optimized data scoring model grounded in SGM to achieve high-quality CT reconstruction for ultra-sparse views. Guan et al.[21] introduced a score-based diffusion model that uses a multi-channel strategy in the projection domain to ensuring that the generated information closer to the original data, resulting in a more accurate SVCT reconstruction. Xia et al.[22] converted the denoising diffusion probability model into a parallel framework to improve the efficiency of the model, and applied it to the reconstruction of breast CT images in dual-domain sparse view. Yang et al.[23] introduced a dual-domain diffusion model for SVCT reconstruction, which includes a sinogram enhancement module and an image refinement module. However, the domain conversion process is limited by the sampling angle, which may lead to artifacts. Although diffusion models have achieved some success in SVCT reconstruction, they still have certain limitations in capturing finer information[24] and rely on large amounts of high-quality data for training[25].
@@
The introduction of mask diffusion mitigates these limitations. Aversa et al.[26] used a layered diffusion model to generate synthetic segmentation masks for high-fidelity diffusion, reducing reliance on labeled data and preserving image detail and structure. Token et al.[27] used a denoising diffusion probability model to simultaneously generate images and masks, addressing data scarcity in satellite segmentation tasks while ensuring a wide diversity of samples. Konz et al.[28] applied mask diffusion to medical image segmentation, demonstrating its advantages when dealing with complex anatomical structures and reducing the reliance on large amounts of high-quality training data.
With the successful application of mask diffusion, it has been gradually introduced into the field of SVCT reconstruction. For example, Tan et al.[29] proposed a score-based multiscale diffusion model. This method introduced a regular mask to reconstruct SVCT, which effectively improved the quality and performance of reconstructed images. However, this limits the network’s ability to learn diverse data distributions, resulting in insufficient model generalization performance.
In order to solve the above issues, we introduce a Sinogram Wavelet Random decomposition And Random mask diffusion Model (SWARM). The random mask strategy effectively expands the limited training sample space, enabling the model to learn a wider range of data distributions. This enhances its understanding of data uncertainty and improves generalization ability. By embedding random masks into sinograms, our method not only ensures that the model captures global information but also enhances its reliability in predicting unfamiliar data distributions. Subsequently, the sinogram wavelet random high-frequency refinement steps further improve the image quality. This strategy improves the model ability to capture details across different frequency bands by enhancing the feature representation, resulting in a reconstructed image with finer texture and clearer detail. By introducing data consistency constraints, the iterative reconstruction process of the two-stage model ensures global information consistency and enhances the accuracy of details such as texture, thereby improving the reliability of the reconstructed image.
@@
$\bullet$ Multi-scale Joint Global-Detail Dual Diffusion. We propose a combined global-detail training method that integrates a sinogram and wavelet high-frequency double diffusion model. By leveraging multi-scale analysis, this approach effectively captures both global information and detailed features in CT reconstruction tasks, thereby significantly enhancing models understanding of image structure.
$\bullet$ Uncertainty Increasing of Random Mask Embedding in Sinogram Training. We propose a training strategy for sinogram with random mask embedding. By randomly generating masks and embedding them into the sinogram, we increase sample diversity and expand the uncertainty of the feature space, thereby improving the model’s generalization ability and enhancing the reliability of the reconstruction results.
$\bullet$ Robust Feature Learning for Sinogram Random High-frequency. The sinogram is decomposed using wavelet transform, and the high-frequency components are randomly selected as network inputs for training. This allows the model to focus on the details of specific frequencies, enhances the representation of spatial features in the data, and improves the model’s adaptability to different frequency features, thereby enhancing its robustness.
@@
2 PRELIMINARY
2.1 Sparse-view CT Image Reconstruction
Given an original image $x$, the image $x$ is converted into full projection data through radon transformation, which is expressed as $R(\theta,s)=\mathfrak{R}(x)(\theta,s)$. $R(\theta,s)$ denotes the outcome of the radon transformation at the position s for angle $\theta$ along the projection direction, and $\mathfrak{R}(x)(\theta,s)$ is the radon transform operator. The projection data obtained through scanning is a linear mapping of noise observation data to a set of measured values, expressed as:
@@ 2.1 Sparse-view CT Image Reconstruction
where $y$ represents the measured projected data, $x$ represents the linear attenuation coefficient distribution of the internal structure of the object. $A$ is the system matrix determined by the geometry of CT equipment and the scanning protocol used. $\varepsilon$ is the system error noise during CT scanning.
SVCT reconstruction is a classical inverse problem [30]. Owing to the incompleteness in the data acquisition process, accurately reconstructing unknown images from limited measurements presents a significant challenge. The mapping from full projection data $y$ to sparse-view data $\hat{y}$ is a process of linear transformation. Specifically, for a linear mapping function $f:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{m\times n}$, acting on $y$ by the linear operator $P(\wedge)$, is obtained and expressed as:
@@ 2.2 Masked Diffusion
Diffusion models have achieved remarkable success in image, audio, and text fields with their high-quality data generation capability. In the forward process, the model gradually transforms the initial data distribution into a Gaussian distribution by progressively adding noise. In the reverse process, the model learns a reverse denoising process to gradually remove the noise, ultimately achieving accurate recovery of the original data. The research on diffusion model mainly focuses on the following aspects:denoising diffusion probabilistic models (DDPMs) [25], score-based generative models (SGMs)[31], and stochastic differential equations(SDEs)[32].
In recent years, masked diffusion has gained widespread application in image processing, particularly in tasks such as image synthesis[33, 34, 35], image editing[36, 37, 38], image restoration[39, 40], image segmentation[41, 42] and image generation[29]. For example, in image synthesis, masks can be used to precisely control the position and shape of generated features, such as in wildfire generation[33]. In image recovery tasks, masks are used to recover occluded or damaged facial areas to achieve the effect of high-fidelity image recovery [39]. Meanwhile, in image segmentation tasks, labeled masks define the target regions, and diffusion models are used to optimize and refine the segmentation results[41]. In addition, in CT image reconstruction tasks, masks are also used in the reconstruction of multi-scale images [29]. These applications show that masks can effectively improve task performance in diffusion models.
Based on the successful application of masks in various image processing tasks, it still falls short when it comes to simulating missing or damaged data in projected data from real scans, we propose an innovative random mask diffusion method for sinogram. By applying random masks to sinograms, our method enables the model to learn from limited data, enhancing its generalization and robustness. Experimental results demonstrate effective improvements in image quality for SVCT reconstruction tasks with masked diffusion.
3 METHOD
3.1 Motivation
In medical imaging, reconstructing high-quality images from sparse projections is vital for accurate diagnosis. When traditional deep models are trained on a closed dataset, satisfactory reconstruction results are achieved within similar data distributions. However, these models tend to decrease diagnostic accuracy and reliability when applied to unfamiliar data, which poses potential clinical risks[43].
Embedding a random mask in a sinogram can enhance the model’s ability to deal with data uncertainty, bolstering its reliability with respect to unfamiliar data distributions. As shown in Fig. 1(a), data distributions with the same or similar data produce relatively uniform results. As shown in Fig. 1(b), a more diverse sample of uncertain data is introduced by embedding random masks in Fig. 1(c), making the data distribution more extensive.
@@ 3.1 Motivation
We provide a theoretical guarantee for the proposed sinogram based random mask embedding strategy and prove the validity of random mask.
Proposition 3.1: The incorporation of random masks into a finite sinogram sample set $x=\{x_{1},x_{2},\cdots,x_{n}\}$, serves to augment the variance of the data distribution.
Proof: Let $x_{M}$ represent the sinogram after random masking $m$. The perturbed data $\tilde{x}$ can be expressed as the sum of the sinogram and its corresponding mask, such that $\tilde{x}=x+x_{M}$. The mean of the perturbed data $\tilde{x}$, denoted as $\tilde{\mu}$, can be formulated as follows:
$\tilde{\mu}=\frac{1}{n}\Sigma^{n}_{i=1}(x_{i}+x_{M_{i}}),$ (4)
@@
$\sigma^{2}(\tilde{x})=\frac{1}{n}\Sigma^{n}_{i=1}((x_{i}-\mu)+(x_{M_{i}}-\mu_{M}))^{2},$ (5)
@@
where $\sigma^{2}(\tilde{x})$ is the variance of $\tilde{x}$, $\mu$ and $\mu_{M}$ represent the mean of $x$ and $x_{M}$, respectively. Depending on the properties of expectation and variance, Eq. (5) can be rewritten as:
$\begin{split}\sigma^{2}(\tilde{x})=\sigma^{2}(x)+\frac{1}{n}\Sigma^{n}_{i=1}2(x_{i}-\mu)(x_{M_{i}}-\mu_{M})\\ +\frac{1}{n}\Sigma^{n}_{i=1}(x_{M_{i}}-\mu_{M})^{2},\end{split}$ (6)
@@
where $\sigma^{2}(x)$ is the variance of $x$. For an arbitrary random mask $m_{i}$, we have $x_{M_{i}}=x_{i}\cdot m_{i}$ and $\mu_{M}=\frac{1}{n}\Sigma^{n}_{j=1}x_{j}\cdot m_{j}$. Extracting the cross terms of Eq. (6), then
$\begin{split}\mathcal{F}&=\frac{1}{n}\Sigma^{n}_{i=1}2(x_{i}-\mu)(x_{M_{i}}-\mu_{M})\\ &=\frac{1}{n}\Sigma^{n}_{i=1}2(x_{i}-\mu)(x_{i}m_{i}-\frac{1}{n}\Sigma^{n}_{j=1}x_{j}m_{j}).\end{split}$ (7)
@@
$\mathcal{F}$ can be decomposed into the sum of two terms, i.e., $\mathcal{F}=\mathcal{F}_{1}-\mathcal{F}_{2}=\frac{1}{n}\Sigma^{n}_{i=1}2(x_{i}-\mu)(x_{i}m_{i})-\frac{1}{n}\Sigma^{n}_{i=1}2(x_{i}-\mu)(\frac{1}{n}\Sigma^{n}_{j=1}x_{j}m_{j})$. It implies that $\mathbb{E}[m_{i}]=\frac{a+b}{2}$ since $m_{i}\sim\mathnormal{U}(a,b)$. Let $\mathbb{E}[\cdot]$ stands for expectation. Then we have $\mathbb{E}[\mathcal{F}_{1}]=\frac{1}{n}\Sigma^{n}_{i=1}2(x_{i}-\mu)x_{i}\mathbb{E}[m_{i}]=\frac{a+b}{2}\cdot\frac{2}{n}\Sigma^{n}_{i=1}(x_{i}-\mu)x_{i}$ =0 since $\Sigma^{n}_{i=1}(x_{i}-\mu)x_{i}=\sigma^{2}(x)$ is the variance term, the expectation is 00. Moreover, $\mathbb{E}[\mathcal{F}_{2}]=-\frac{2}{n^{2}}\Sigma^{n}_{i=1}\Sigma^{n}_{j=1}(x_{i}-\mu)\cdot x_{j}\cdot\frac{a+b}{2}=0$ since $\Sigma^{n}_{i=1}(x_{i}-\mu)=0$. Hence, $\mathbb{E}[\mathcal{F}]=0$.
@@
Since $\mathbb{E}[\frac{1}{n}\Sigma^{n}_{i=1}(x_{M_{i}}-\mu_{M})^{2}]\geq 0$, it means that $\mathbb{E}[\sigma^{2}(\tilde{x})]\geq\sigma^{2}(x)$. Therefore, the random mask strategy we introduce to the data set is effective. $\square$
The high-frequency components extracted by wavelet transform are essential for capturing complex texture details in images. The method of randomly embedding high-frequency components can enhance the ability of the model to recognize and capture image details, so as to improve the performance of the model in the reconstruction of detail information. This method helps to improve the stability and accuracy of the model, ensuring that the denoised image can truly reflect the fine features of the original scene.
Through two-stage iterative optimization process, the algorithm ensures global consistency in images while generating high-quality ones with rich texture and clear details. This method effectively enhances the accuracy and robustness of the reconstruction results, ensuring the fidelity and precision of the sinogram, and achieving more accurate and reliable image reconstruction.
@@
Figure 2: Training stage: (a) A model training based on random masks in sinogram. (b) A model training for high-frequency random decomposition of wavelet based on sinogram. Iteration reconstruction stage: (c) The proposed SWARM method is used to reconstruct the sparse-view CT projection domain. “LF”: Low-frequency. “HF”: High-frequency.
3.2 Training Process in Sinogram Wavelet Domain
3.2.1 Random Mask Strategy in Sinogram Domain

### v2

@@ Abstract
Diffusion models have demonstrated strong potential in sparse-view computed tomography (SVCT) reconstruction. However, their generalization ability is often constrained when trained on limited sample spaces, leading to performance degradation when encountering unseen data. This typically leads to image blurring, loss of structural details, and cross-region inconsistencies. To address these challenges, we propose a Sinogram-based Wavelet random decomposition And Random mask diffusion Model (SWARM) for SVCT reconstruction. Specifically, introducing a random mask strategy in the sinogram effectively expands the limited training sample space. This enables the model to learn a broader range of data distributions, enhancing its understanding and generalization of data uncertainty. In addition, we designed a wavelet-based random training mechanism for sinogram high-frequency components, enabling the model to capture structural details in different frequency bands and enhancing the richness and structural consistency of the representations. Two-stage iterative reconstruction method is adopted to ensure the global consistency of the reconstructed image while refining the details. Compared with other state-of-the-art reconstruction methods, SWARM can increase the PSNR by 3.59 dB on average, the SSIM by 0.69%, and reduce the MSE by 55.40%. These experimental results indicate that SWARM has great potential in the field of sparse-view CT image reconstruction.
Index Terms:
@@ I Introduction
In recent years, diffusion models have been increasingly employed in SVCT reconstruction owing to their superior capabilities. Xia et al.[19] introduced a patch-based denoising probabilistic diffusion model to improve SVCT reconstruction performance and address large memory requirements. Additionally, Wu et al.[20] proposed an iteratively optimized data scoring model grounded in SGM to achieve high-quality CT reconstruction for ultra-sparse views. Guan et al.[21] introduced a score-based diffusion model that uses a multi-channel strategy in the projection domain to ensuring that the generated information closer to the original data, resulting in a more accurate SVCT reconstruction. Xu et al.[22] used a diffusion model for stage-by-stage optimization in the wavelet domain, which enhanced sparse-view CT image quality. Xia et al.[23] converted the denoising diffusion probability model into a parallel framework to improve the efficiency of the model, and applied it to the reconstruction of breast CT images in dual-domain sparse view. Yang et al.[24] introduced a dual-domain diffusion model for SVCT reconstruction, which includes a sinogram enhancement module and an image refinement module. Although diffusion models have achieved some success in SVCT reconstruction, they still have certain limitations in capturing finer information[25] and rely on large amounts of high-quality data for training[26].
@@
To partially address these challenges, masked diffusion encourages the model to attend to critical regions or structural features, thereby enhancing its ability to reconstruct local details in data-scarce settings. Moreover, introducing perturbations into the input data can enhance the robustness of the model against missing information and improve the generalization performance of the model. Aversa et al.[27] used a layered diffusion model to generate synthetic segmentation masks for high-fidelity diffusion, reducing reliance on labeled data and preserving image detail and structure. Toker et al.[28] used a denoising diffusion probability model to simultaneously generate images and masks, addressing data scarcity in satellite segmentation tasks while ensuring a wide diversity of samples. Konz et al.[29] applied masked diffusion to medical image segmentation, demonstrating its advantages when dealing with complex anatomical structures and reducing the reliance on large amounts of high-quality training data. With the successful application of masked diffusion, it has been gradually introduced into the field of SVCT reconstruction. For example, Tan et al.[30] proposed a score-based multiscale diffusion model. By introducing a regular mask into the SVCT reconstruction framework, this approach effectively improved model performance. Nevertheless, the reliance on fixed masking strategies imposes constraints on the network’s capacity to learn diverse data distributions, thereby limiting the generalization of the model.
To address the above issues, we introduce a Sinogram Wavelet random decomposition And Random mask diffusion Model (SWARM). The proposed random mask strategy effectively expands the limited training sample space, allowing the model to learn from a broader range of data distributions. This dual mechanism not only reinforces the model’s capacity to characterize epistemic uncertainty but also substantially elevates its generalization capability. By integrating stochastic masking into sinogram projections, our approach fosters a synergistic learning framework that enables the model to encode global structural dependencies while enhancing its robustness in extrapolating to unseen data distributions. To further refine reconstruction quality, we apply wavelet-based random high-frequency decomposition, which improves the representation of fine structural details across frequency bands. Finally, a two-stage iterative reconstruction framework with data consistency constraints ensures both global structural coherence and accurate local detail recovery.
@@
$\bullet$ Multi-scale Joint Global-Detail Dual Diffusion. We propose a global-detail integrated training framework that incorporates a dual diffusion model. By harnessing multi-scale analysis, this methodology effectively captures both global contextual information and fine-grained features in CT reconstruction tasks, thereby significantly enhancing the model’s comprehension of image structural hierarchies.
$\bullet$ Uncertainty Increasing of Random Mask Embedding in Sinogram Training. We introduce a strategy of embedding sinogram with random masks to simulate incomplete sampling scenarios, thereby augmenting uncertainty and diversity in the training process. This approach induces feature-space uncertainty, enriches training data diversity, and enhances model generalization capability as well as reconstruction robustness under distribution shifts.
$\bullet$ Robust Feature Learning for Sinogram Random High-frequency. The sinogram is decomposed into multi-frequency subbands via wavelet transform, with high-frequency components undergoing random sampling to serve as training inputs. This frequency-sensitive sampling strategy enhances spatial feature discriminability and fortifies the model’s robustness against structural variations and noise perturbations.
@@
II Preliminary
II-A Sparse-view CT Image Reconstruction
Given an original image $x$, it represents the linear attenuation coefficient distribution of internal structures within the object, reflecting the differential absorption characteristics of heterogeneous tissues to X-ray photons. The image $x$ is converted into full projection data through radon transformation, which is expressed as $R(\theta,s)=\mathfrak{R}(x)(\theta,s)$. $R(\theta,s)$ denotes the outcome of the radon transformation at the position s for angle $\theta$ along the projection direction. $\mathfrak{R}(x)(\theta,s)$ is the radon transform operator. The projection data obtained through scanning is a linear mapping of noise observation data to a set of measured values, expressed as:
@@ II-A Sparse-view CT Image Reconstruction
where $y$ represents the measured projected data, $A$ is the system matrix determined by the geometry of CT equipment and the scanning protocol used and $\varepsilon$ represents the system error and random noise. SVCT reconstruction is a classical inverse problem [31]. Owing to the incompleteness in the data acquisition process, accurately reconstructing unknown images from limited measurements presents a significant challenge. The mapping from full projection data $y$ to sparse-view projection data $\hat{y}$ is a process of linear transformation. Specifically, for a linear mapping function $f:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{m\times n}$, acting on $y$ by the linear operator $P(\wedge)$, is obtained and expressed as:
@@ II-B Masked Diffusion
Diffusion models have achieved remarkable success in image, audio, and text fields with their high-quality data generation capability. In the forward process, the model gradually transforms the initial data distribution into a gaussian distribution by progressively adding noise[32, 33, 34, 35]. In the reverse process, the model learns a reverse denoising process to gradually remove the noise, ultimately achieving accurate recovery of the original data. The research on diffusion model mainly focuses on the following aspects: denoising diffusion probabilistic models (DDPMs) [26], score-based generative models (SGMs)[36], and stochastic differential equations(SDEs)[37].
Masked diffusion has recently emerged as a powerful technique in image processing, demonstrating strong capabilities in learning data representations from incomplete inputs. By applying binary masks to input images, masked diffusion enables models to reconstruct missing regions, thus enforcing semantic understanding and structural reasoning. This masking strategy encourages the model to focus on global coherence while generating local details, which has proven effective in self-supervised and unsupervised settings alike.
While existing works have leveraged masked diffusion in diverse tasks such as image synthesis[38, 39, 27], image editing[40, 41, 42], image restoration[43, 44], image segmentation[45, 46]. Most approaches focus on natural images and overlook the unique challenges of projection data in medical imaging. In particular, conventional masking strategies fail to simulate the complex patterns of missing or corrupted sinogram data encountered in SVCT reconstruction. To address this gap, by introducing random occlusion in the projection domain, our method enhances the model’s ability to generalize from finite and degenerate inputs.
III Method
III-A Motivation
In medical imaging, reconstructing high-quality images from sparse-view projections is vital for accurate diagnosis. Although traditional deep learning models achieve satisfactory performance when trained and tested within closed, homogeneous datasets. Their generalization ability degrades significantly when deployed on out-of-distribution (OOD) or clinically diverse data. Abdar et al. [47] and Huang et al. [48] emphasize the risk of overfitting to training distributions and the consequent drop in diagnostic reliability. However, current literature rarely explores effective strategies to explicitly enhance robustness to distribution shifts in the projection domain, especially in the context of SVCT. Most approaches focus on either improving reconstruction architectures or increasing data diversity through simple augmentation, failing to systematically address the sensitivity of models to the variability in sampling patterns and acquisition conditions.
To address this gap, we propose a novel training strategy that embeds structured randomness into the projection data via random masks. By perturbing full-view sinograms during training, we simulate diverse acquisition scenarios, encouraging the model to learn robust, generalizable reconstruction mappings. As illustrated in Fig. 1(b)-(c), the embedded random masks induce broader and more uncertain data distributions, which can improve the model’s robustness to unseen clinical variations.
@@ III-A Motivation
To theoretically understand how perturbed data influences data diversity generation mechanisms, this paper constructs a theoretical interpretive framework and systematically analyzes full-view projection data and the random masking of such data.
Proposition 3.1: The incorporation of random masks into a finite sinogram sample set $y=\{y_{1},y_{2},\cdots,y_{n}\}$, serves to augment the variance of the data distribution.
Proof: Let $y_{M}$ represent the sinogram after random masking $m$. The perturbed data $\tilde{y}$ can be expressed as the sum of the sinogram and its corresponding mask, such that $\tilde{y}=y+y_{M}$. The mean of the perturbed data $\tilde{y}$, denoted as $\tilde{\mu}$, can be formulated as follows:
$\tilde{\mu}=\frac{1}{n}\Sigma^{n}_{i=1}(y_{i}+y_{M_{i}}).$ (4)
@@
$\sigma^{2}(\tilde{y})=\frac{1}{n}\Sigma^{n}_{i=1}((y_{i}-\mu)+(y_{M_{i}}-\mu_{M}))^{2},$ (5)
@@
where $\sigma^{2}(\tilde{y})$ is the variance of $\tilde{y}$, $\mu$ and $\mu_{M}$ represent the mean of $y$ and $y_{M}$, respectively. Depending on the properties of expectation and variance, Eq. (5) can be rewritten as:
$\begin{split}\sigma^{2}(\tilde{y})=\sigma^{2}(y)+\frac{1}{n}\Sigma^{n}_{i=1}2(y_{i}-\mu)(y_{M_{i}}-\mu_{M})\\ +\frac{1}{n}\Sigma^{n}_{i=1}(y_{M_{i}}-\mu_{M})^{2},\end{split}$ (6)
@@
where $\sigma^{2}(y)$ is the variance of $y$. For an arbitrary random mask $m_{i}$, we have $y_{M_{i}}=y_{i}\cdot m_{i}$ and $\mu_{M}=\frac{1}{n}\Sigma^{n}_{j=1}y_{j}\cdot m_{j}$. Extracting the cross terms of Eq. (6), then
$\begin{split}\mathcal{F}&=\frac{1}{n}\Sigma^{n}_{i=1}2(y_{i}-\mu)(y_{M_{i}}-\mu_{M})\\ &=\frac{1}{n}\Sigma^{n}_{i=1}2(y_{i}-\mu)(y_{i}m_{i}-\frac{1}{n}\Sigma^{n}_{j=1}y_{j}m_{j}).\end{split}$ (7)
@@
$\mathcal{F}$ can be decomposed into the sum of two terms, i.e., $\mathcal{F}=\mathcal{F}_{1}-\mathcal{F}_{2}=\frac{1}{n}\Sigma^{n}_{i=1}2(y_{i}-\mu)(y_{i}m_{i})-\frac{1}{n}\Sigma^{n}_{i=1}2(y_{i}-\mu)(\frac{1}{n}\Sigma^{n}_{j=1}y_{j}m_{j})$. It implies that $\mathbb{E}[m_{i}]=\frac{a+b}{2}$ since $m_{i}\sim\mathnormal{U}(a,b)$. Let $\mathbb{E}[\cdot]$ stands for expectation. Then we have $\mathbb{E}[\mathcal{F}_{1}]=\frac{1}{n}\Sigma^{n}_{i=1}2(y_{i}-\mu)y_{i}\mathbb{E}[m_{i}]=\frac{a+b}{2}\cdot\frac{2}{n}\Sigma^{n}_{i=1}(y_{i}-\mu)y_{i}$ =0 since $\Sigma^{n}_{i=1}(y_{i}-\mu)y_{i}=\sigma^{2}(y)$ is the variance term, the expectation is 00. Moreover, $\mathbb{E}[\mathcal{F}_{2}]=-\frac{2}{n^{2}}\Sigma^{n}_{i=1}\Sigma^{n}_{j=1}(y_{i}-\mu)\cdot y_{j}\cdot\frac{a+b}{2}=0$ since $\Sigma^{n}_{i=1}(y_{i}-\mu)=0$. Hence, $\mathbb{E}[\mathcal{F}]=0$.
@@
Since $\mathbb{E}[\frac{1}{n}\Sigma^{n}_{i=1}(y_{M_{i}}-\mu_{M})^{2}]\geq 0$, it means that $\mathbb{E}[\sigma^{2}(\tilde{x})]\geq\sigma^{2}(x)$. $\square$
The derivation shows that perturbing the data increases data variance, thereby expanding the distribution from which the model learns. This heuristic improvement in generalization arises from exposing the model to a wider range of data. Notably, the benefit of increased variance hinges on the compatibility between the perturbed data distribution and the target test distribution. Thus, our variance-based formulation should be viewed as a heuristic perspective. Our subsequent experiments (Section IV) empirically confirm that this controlled stochasticity enhances generalization across SVCT reconstruction tasks.
To further enhance image reconstruction fidelity, we leverage the high-frequency components extracted by wavelet transforms, which are crucial for preserving fine structures such as edges and textures. During training, we introduce a randomized perturbation strategy applied specifically to high-frequency wavelet subbands. This enhances the model’s sensitivity to microstructural details and improves its capability to represent multi-scale high-frequency information, thereby boosting detail recovery and robustness under noisey and incomplete data conditions. By integrating high-frequency priors with inherent uncertainty, the model achieves improved generalization while maintaining structural fidelity. Additionally, a two-stage iterative optimization framework is employed to jointly ensure global structural coherence and accurate high-frequency detail reconstruction. This dual-phase strategy balances fidelity and texture synthesis, enabling robust and high-quality reconstructions across diverse imaging scenarios.
@@
Figure 2: The pipeline of SWARM training process and iterative reconstruction procedure. Training stage: (a) A model training based on random masks in sinogram. (b) A model training for high-frequency random decomposition of wavelet based on sinogram. Iteration reconstruction stage: (c) The proposed SWARM method is used to reconstruct the sparse-view CT projection domain. “LF”: Low-frequency. “HF”: High-frequency.
III-B Training Process in Sinogram Wavelet Domain